OpenStudy (kayders1997):

Please help!!!!!! Just want to make sure I'm doing this linearization right f(x)=x^3 a=2 and g(x)=(x)^1/3

OpenStudy (kayders1997):

A=-27

OpenStudy (kayders1997):

@zepdrix @Directrix

OpenStudy (kayders1997):

@mathmale

OpenStudy (kayders1997):

Wait sorry 12x-16

Directrix (directrix):

Is this what you are doing? See attachment.

OpenStudy (kayders1997):

Yes

Directrix (directrix):

I can't read the question. >> f(x)=x^3 a=2 Was something left out?

OpenStudy (kayders1997):

Oh just find the linearization L(x) of the function at a

OpenStudy (kayders1997):

Don't know @Directrix it's okay if you don't :)

OpenStudy (kayders1997):

I know you have to find the derivative

Directrix (directrix):

Look at this attachment and see what you think. http://www.emathhelp.net/calculators/calculus-1/linear-approximation-calculator/

Directrix (directrix):

Is this a second problem for linearization --> g(x)=(x)^1/3

OpenStudy (kayders1997):

Yes

OpenStudy (kayders1997):

Thank you for the website!!!! :)

OpenStudy (kayders1997):

@RhondaSommer

OpenStudy (kayders1997):

The last one won't work at @Directrix! :(

OpenStudy (kayders1997):

I'm stuck on second one

RhondaSommer (rhondasommer):

hmmm hold on

OpenStudy (kayders1997):

Okay

Directrix (directrix):

You have to read the instructions on how to enter the function. @kayders1997